A heating element is made by maintaining a potential difference of 66.0 V across the length of a Nichrome wire that has a 3.15 × 10–6 m2 cross section. Nichrome has a resistivity of 5.00 × 10–7 Ω·m. (a) If the element dissipates 4910 W, what is its length? (b) If 113 V is used to obtain the same dissipation rate, what should the length be?
Answer #1

The power dissipated is P= i^2R =V^2/R

where R = ρL A/ , ρ is the resistivity, L is the length and A = πr 2 is the cross sectional area.

R = 5 x 10^-7 x L / 3.15 x 10^-6 = .1587L ohm

P = 4910 watt & V = 66 Volt
=>By P = V^2/R
=>4910 = 66^2/.1587L
=>L = 5.59m
(b) =>4910 = 113^2/.1587L

L = 16.387m